2 points Beyoncáo has nine background dancers for her performance at the Grammys, three of them are wearing black costumes and six are wearing white costumes. If two of the dancers are selected at random to join BeyoncAD onstage for the encore, what is the probability that at least one will be wearing black given that at least one is wearing white? Enter your answer as a traction in lowest terms
Answer #1

We need to find the probability that one of the dancers selected will be wearing black costume if one is already wearing white.

There are 3 black costumes and 5 white costumes (one costume is already white)

So probability that one black will be selected = 3/(3+5) = 3/8